OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (-2, 1). Which equation defines circle P?

OpenStudy (anonymous):

(x + 2)2 + (y – 1)2 = 8 (x – 2)2 + (y + 1)2 = 8 (x + 2)2 + (y – 1)2 = 64 (x – 2)2 + (y + 1)2 = 64

OpenStudy (anonymous):

the standard equation of a circle with center (h, k) and radius r is given as: \[\huge (x-h)^2+(y-k)^2=r^2 \] which one of those choices will fit this equation?

OpenStudy (anonymous):

idk ? im not sure how to do this D:

OpenStudy (anonymous):

what's the (h, k) , the center, in your problem? give me the coordinates...

OpenStudy (anonymous):

-2 & 1

OpenStudy (anonymous):

so h=-2, k=1.. can you plug that into the equation i gave u?

OpenStudy (anonymous):

(x-2)^2 + ( y-1)^2=8^2 ?

OpenStudy (anonymous):

but i thought you said h=-2,, so shouldn't it start as (x - -2)^2 ?

OpenStudy (anonymous):

oops, i thought i hit the - button twice. but yes it should be (x--2)^2

OpenStudy (anonymous):

ok... without giving me the choices, just replace the given values into the equation i gave....

OpenStudy (anonymous):

after simplifying it, which is the correct answer?

OpenStudy (anonymous):

(x + 2)2 + (y – 1)2 = 64 ?

OpenStudy (anonymous):

that's the one....:)

OpenStudy (anonymous):

okay ' now it makes sense. can you explain how to find the center and radius of this circle? please .

OpenStudy (anonymous):

from a graph

OpenStudy (anonymous):

ok... close shop here and post it as another question....:)

OpenStudy (anonymous):

okay'

OpenStudy (anonymous):

ok...:)